### v1

# Revisiting the Many Instruments Problem using Random Matrix Theory

Abstract
We use recent results from the theory of random matrices to improve instrumental variables estimation with many instruments. In settings where the first-stage parameters are dense, we show that Ridge lowers the implicit price of a bias adjustment. This comes along with improved (finite-sample) properties in the second stage regression. Our theoretical results nest existing results on bias approximation and bias adjustment. Moreover, it extends them to settings with more instruments than observations.
1 Introduction
In this paper we propose a new bias-adjusted Ridge estimator for a simultaneous equations model with a single endogenous regressor and many instrumental variables (IVs). Our motivation comes from recent results in the theory of random matrices (RMT), which give a new understanding of the bias induced by many instruments.
IV estimators perform poor under many and/or weak instruments, which has been extensively discussed in the literature (e.g., Bound et al., 1995; Staiger and Stock, 1997; Stock and Wright, 2000; Chao and Swanson, 2005). The standard IV estimator is a two-stage least squares (2SLS) procedure, where we first use least squares to predict the endogenous variable, and in a second least squares regression use the predicted values to estimate the causal effect. Thus, the accuracy of 2SLS estimation depends crucially on the quality of the prediction in the first-stage regression. We would expect a method that reduces the predictive risk in the first stage to perform better in the second-stage regression. This reasoning motivated a series of papers to replace OLS in the first stage by statistical learning tools (such as Belloni et al. (2012), Gold et al. (2020) and Angrist and Frandsen (2022) for Lasso regressions or Okui (2011) and Hansen and Kozbur (2014) for Ridge regressions). In this paper, we are interested in settings where Ridge is the preferred choice in the first-stage regression.
The goal of this paper is to gain a better understanding of how a first-stage Ridge regression affects the estimation of the causal effect in a high-dimensional setting. We show how RMT can be used to adjust for the many IV bias. We first consider a standard many IV setting where the number of instruments is large relative to, but still smaller than the sample size. In this setting, our bias approximation nests the well-known results for least squares in the first stage (e.g., Hahn and Hausman, 2002a, b; Bun and Windmeijer, 2011), and a special case of our bias adjustment coincides with the well-known bias-adjusted 2SLS estimator (Nagar, 1959; Chao and Swanson, 2005; Donald and Newey, 2001). Bias adjustments are costly in the sense that they reduce the signal strength of the instruments. We show that a first-stage Ridge regression reduces the implicit price for a bias adjustment, which comes along with improved (finite-sample) properties in the second stage. Finally, the results mentioned above generalize to settings where the number of instruments is larger than the number of observations, which has not been discussed in the IV literature so far.
To derive our results, we make the following assumptions. We assume that each of the instruments has only a small, independent random effect on the endogenous variable. This can be considered an average-case analysis over dense parameters. Chamberlain and Imbens (2004) also use a random effects assumption to derive a quasi-maximum likelihood estimator in an instrumental variables regression with many instruments. Moreover, we assume a high dimensional asymptotic regime for the instrumental variables. In such a setting we would expect Ridge to outperform OLS with respect to prediction risk (e.g., Dobriban and Wager, 2018a, among others). These assumptions, while being standard in the RMT literature, are, of course, very restrictive. However, imposing them allows us to derive insightful results that nest existing approaches, and may motivate future conceptual developments.
Closest to our study in the RMT literature is Dobriban and Wager (2018a) and Hastie et al. (2022b) who analyze the performance of Ridge in a single equation model. In the econometrics literature our study is close to Hahn and Hausman (2002a, b) and Bun and Windmeijer (2011) which analyze the bias of 2SLS under many and/or weak instruments and, additionally, close to several econometric approaches that adjust for the 2SLS bias (e.g., Nagar, 1959; Chamberlain and Imbens, 2004; Chao and Swanson, 2005; Hansen and Kozbur, 2014). In economics our study is closely related to Angrist and Frandsen (2022) who compare bias-adjusted IV estimators with estimators based on Lasso in the first stage regression. They conclude that statistical learning techniques seem to be less encouraging in IV settings. In contrast to this conclusion, we show that bias-adjusted 2SLS estimators can indeed be motivated by results from the statistical learning literature. This allows us to improve the bias adjustment, which leads to a better finite-sample performance. Additionally, it suggests that a comparison between bias-adjusted IV estimators and estimators based on Lasso essentially boils down to the distinction between dense or sparse first-stage parameters.
2 Model and 2SLS Estimation
We have a random sample $\left\{y_{i},x_{i},z_{i}^{\prime}\right\}_{i=1}^{n}$ drawn from the following simultaneous equations model
$y_{i}$ $=$ $\beta x_{i}+\varepsilon_{i}$ (2.1)
$x_{i}$ $=$ $z_{i}^{\prime}\pi+\nu_{i}$ (2.2)
where $y_{i}$ and $x_{i}$ are the entries of the $n\times 1$ vectors $y$ and $x$, and $z_{i}$ are arranged as the rows of the $n\times k$ matrix $Z$. The aim is to estimate the causal effect $\beta$. We assume throughout that the following model conditions hold.
Assumption 1 (Model).
Let $\eta_{i}=\left(\varepsilon_{i}\leavevmode\nobreak\ \nu_{i}\right)^{\prime}$ be normally distributed independent of the other random quantities with $E\left[\eta_{i}\right]=0$ and $E\left[\eta_{i}\eta_{i}^{\prime}\right]=\left[\begin{array}[]{cc}\sigma_{\varepsilon}^{2}&\sigma_{\varepsilon\nu}\\ \sigma_{\varepsilon\nu}&\sigma_{\nu}^{2}\end{array}\right]$, a positive definite matrix.
In particular, Assumption 1 imposes homoskedasticity and joint normality of $\epsilon$ and $\nu$. If $\sigma_{\varepsilon\nu}\neq 0$, $x$ is endogenous and the OLS estimate of $\beta$ is inconsistent. The instrumental variables $Z$ can be used to overcome the endogeneity. The standard IV estimator is a two-stage least squares procedure (2SLS). First, we use least squares to predict $\widehat{x}=P_{0}x$ with $P_{0}=Z(Z^{\prime}Z)^{-1}Z^{\prime}\,$. In a second least squares regression we use $\widehat{x}\,$ to estimate the causal effect $\beta$. A first-stage prediction based on Ridge is defined as $\widehat{x}_{\lambda}=P_{\lambda}x$ with $P_{\lambda}=Z(Z^{\prime}Z/n+\lambda I_{k})^{-1}Z^{\prime}/n\,$ being the Ridge smoother and $\lambda$ being a tuning parameter. Clearly, when we set $\lambda\downarrow 0$ in settings with less instruments than observations, we get 2SLS. If we set $\lambda\downarrow 0$ in settings with more instruments than observations, we get a ridgeless 2SLS estimator. For a detailed discussion of ridgeless least squares and its interpretation as min-norm least squares estimator, see for example Hastie et al. (2022b).
Following the RMT literature, we consider a high-dimensional asymptotic regime for the instrumental variables.
Assumption 2 (High-Dimensional Asymptotics).
The following conditions hold:
  • a)
    The data $Z\in\mathbb{R}^{n\times k}$ are generated as $Z=W\Sigma^{1/2}$ for an $n\times k$ matrix $W$ with i.i.d. entries satisfying $\textrm{E}[W_{ij}]=0$ and $\textrm{Var}[W_{ij}]=1$, and a deterministic $k\times k$ positive semidefinite covariance matrix $\Sigma\,$.
  • b)
    The sample size $n\to\infty$ and the number of instruments $k\to\infty$, such that the ratio $k/n\to\gamma>0\,$.
  • c)
    The spectral distribution $F_{\Sigma}$ of $\Sigma$ converges to a limit probability distribution $H$ supported on $[0,\infty)$, called the population spectral distribution (PSD).
See Dobriban and Wager (2018a) for a detailed discussion of covariance matrices $\Sigma$ that are aligned with this setting. For example, the autoregressive AR-1 model with $\Sigma_{ij}=\rho^{|i-j|}$, which is often used in econometrics, fits Assumption 2.
Following the RMT literature, we, moreover, assume that each of the instruments has only a small and independent random effect on the endogenous variable.
Assumption 3 (Random Coefficients).
The first-stage parameters $\pi\in\mathbb{R}^{k}$ are random with $\textrm{E}[\pi]=0\,$, $\textrm{Var}[\pi]=k^{-1}\alpha^{2}I_{k}$ and $\alpha^{2}\neq 0\,$.
This assumption has been used before in the many instruments literature by Chamberlain and Imbens (2004) in a theoretical analysis without RMT. The concentration parameter, $\mu^{2}\equiv n\alpha^{2}/\sigma_{\nu}^{2}\,$, is a common measure of instrument strength. In a weak instrument setting the concentration parameter is small relative to the sample size.
Throughout the paper we illustrate our theoretical results using Monte Carlo simulations based on a standard many IV setting
$y$ $=$ $\beta x+\varepsilon$
$x$ $=$ $Z\pi+\nu$
with $\varepsilon=\rho\nu+\sqrt{1-\rho^{2}}w$ where $\nu\sim N(0,1)$ and $w\sim N(0,1)$ are independent ($\rho=0.6$). Moreover, $\pi\sim N(0,\sigma^{2}_{\pi})$ with $\sigma^{2}_{\pi}=\alpha^{2}/k$ and $\alpha^{2}=\gamma F\,$. This allows us to control the first-stage $F$ statistic, which we generally set equal to 5. We show simulation results for $Z\sim N(0,\Sigma)$ and a) isotropic $\Sigma=I_{k}$ or b) AR-1 model $\Sigma_{ij}=0.5^{|i-j|}$. Additionally, we set $n=200$, $k=150$ or $k=250$, and $\beta=0$.
3 Results from Random Matrix Theory
We start with a brief review of key results from the theory of random matrices, which help us to describe and adjust for the bias induced by Ridge estimation in the first-stage regression. From Marchenko and Pastur (1967) and Silverstein (1995) we know that the spectral distribution of the sample covariance matrix $\widehat{\Sigma}=Z^{\prime}Z/n$ converges weakly, with probability 1, to a limiting distribution. This limiting distribution is called the empirical spectral distribution (ESD). Let $Y$ be a random variable distributed according to the ESD of $\widehat{\Sigma}$, then the well-known Stieltjes transform $m(-\lambda)$ is
$m(-\lambda)=\textrm{E}\left[\frac{1}{Y+\lambda}\right].$
The Stieltjes transform of the spectral measure of $\widehat{\Sigma}$ satisfies
$\widehat{m}(-\lambda)=\frac{1}{k}\operatorname{tr}\big{(}(Z^{\prime}Z/n+\lambda I_{k})^{-1}\big{)}\to_{a.s.}m(-\lambda)\,.$
Additionally, we need the companion Stieltjes transform $v(-\lambda)$, which is the Stieltjes transform of the ESD of the $n\times n$ matrix $\underline{\widehat{\Sigma}}=ZZ^{\prime}/n\,$. Let $\underline{Y}$ be a random variable distributed according to the ESD of $\underline{\widehat{\Sigma}}\,$, then
$v(-\lambda)=\textrm{E}\left[\frac{1}{\underline{Y}+\lambda}\right],$
and the Stieltjes transform of the spectral measure of $\underline{\widehat{\Sigma}}$ satisfies
$\widehat{v}(-\lambda)=\frac{1}{n}\operatorname{tr}\big{(}(ZZ^{\prime}/n+\lambda I_{n})^{-1}\big{)}\to_{a.s.}v(-\lambda)\,.$
To calculate the population quantities $m(-\lambda)$ and $v(-\lambda)$ in the simulations, we use the procedure described in Dobriban and Wager (2018b). To estimate them in a sample, we follow Dobriban and Sheng (2020) and use $\widehat{m}(-\lambda)$ and $\widehat{v}(-\lambda)\,$.
In the following, we make repeatedly use of well-known results in the RMT literature (the proofs are replicated in Appendix A for convenience). First, we mention a useful representation of the Ridge smoother.
Lemma 3.1.
$P_{\lambda}=I_{n}-\lambda(ZZ^{\prime}/n+\lambda I_{n})^{-1}\,.$
An interesting implication of this is that $M_{\lambda}=I_{n}-P_{\lambda}=\lambda(ZZ^{\prime}/n+\lambda I_{n})^{-1}\,$. Second, we mention a useful link between the Stieltjes transform and its companion.
Lemma 3.2.
$\lambda v(-\lambda)=1-\gamma(1-\lambda m(-\lambda))\,.$
Third, we mention limits involving the Stieltjes transforms that help us to work out the links to the existing literature on many weak instruments.
Lemma 3.3.
For the ridgeless limit ($\lambda\downarrow 0$), we have
  • if $\gamma<1$, then  1a) $\lim_{\lambda\downarrow 0}\lambda v(-\lambda)=1-\gamma$  and  1b) $\lim_{\lambda\downarrow 0}\lambda m(-\lambda)=0\,,$
  • if $\gamma>1$, then  2a) $\lim_{\lambda\downarrow 0}\lambda v(-\lambda)=0$  and  2b) $\lim_{\lambda\downarrow 0}\lambda m(-\lambda)=1-\frac{1}{\gamma}\,.$
We, moreover, make use of the Silverstein equation (Silverstein, 1995), which is
[…]
where $\frac{dH(t)}{dt}$ is the pdf of the population spectral distribution. The Silverstein equation links the limit PSD to the limit ESD. If instruments are equi-correlated, $\Sigma$ has $k-1$ eigenvalues that equal $1-\rho_{z}$, so the pdf of the PSD is a point mass at $1-\rho_{z}$. Therefore, the Silverstein equation has a closed-form expression
$\frac{1}{v(-\lambda)}=\lambda+\gamma\frac{1-\rho_{z}}{1+(1-\rho_{z})v(-\lambda)}\,.$
In the special case of isotropic instruments ($\rho_{z}=0$), this simplifies to
$\frac{1}{v(-\lambda)}=\lambda+\frac{\gamma}{1+v(-\lambda)}$
(see also Hastie et al., 2022a).
4 2SLS Estimation and RMT
4.1 Bias Approximation
We derive a bias approximation for the causal effect $\beta$ when the first-stage regression is based on Ridge. Consider the first-stage Ridge estimates $\widehat{\pi}_{\lambda}$ from a regression of the endogenous variable $x$ on the instruments $Z$ given by
$\widehat{\pi}_{\lambda}=\operatorname*{arg\,min}_{\pi}\left\{\frac{1}{n}\|x-Z\pi\|^{2}_{2}+\lambda\|\pi\|^{2}_{2}\right\}\,,$
or, equivalently written as $\widehat{\pi}_{\lambda}=\big{(}Z^{\prime}Z/n+\lambda I_{k}\big{)}^{-1}Z^{\prime}x/n\,$. In the second stage we use the predicted values $\widehat{x}_{\lambda}=Z\widehat{\pi}_{\lambda}$ in a just-identified 2SLS regression to obtain
$\widetilde{\beta}_{\lambda}=\frac{x^{\prime}P_{\lambda}y}{x^{\prime}P_{\lambda}x}\,.$ (4.1)
In the following we call $\widetilde{\beta}_{\lambda}$ the 2SLS-Ridge estimator to emphasize that it is a two-stage procedure with Ridge in the first stage and just-identified 2SLS in the second stage regression. For the limit $\lambda\downarrow 0$ we get the (ridgeless) 2SLS estimator, which corresponds to standard 2SLS if $\gamma<1$. As already mentioned in Hansen and Kozbur (2014), 2SLS-Ridge is biased. Theorem 4.1 gives precise information about the bias induced by using Ridge in the first-stage regression. While the numerator only contains a noise term, the denominator is the sum of a signal term and a noise term.
Theorem 4.1.
Under Assumptions 1, 2 and 3, suppose additionally that the eigenvalues of $\Sigma$ are uniformly bounded away from zero and infinity, and $\textrm{E}[W_{ij}^{8}]$ are uniformly bounded from above. Then the bias of 2SLS-Ridge estimation converges almost surely
$\widetilde{\beta}_{\lambda}-\beta\to_{a.s.}\frac{\sigma_{\varepsilon\nu}\big{(}1-\lambda v(-\lambda)\big{)}}{\alpha^{2}\big{(}1-\lambda p(-\lambda)\big{)}+\sigma_{\nu}^{2}(1-\lambda v(-\lambda))}$
with $p(-\lambda)\equiv 1-\lambda m(-\lambda)$.
Proof. See Appendix B. ∎
Before discussing ways to get rid of the bias term later in Section 4.2, we first briefly address the links of our bias approximation to the existing literature and the advantage of using Ridge compared to least squares in the first-stage regression. The first part of the next corollary replicates the well-established 2SLS bias approximation using Theorem 4.1 and Lemma 3.3 (compare, for example, Hahn and Hausman (2002a, b) and Chao and Swanson (2005)). The second part of the corollary extends these results to ridgeless 2SLS estimation.
Corollary 4.1.1.
For the ridgeless limit ($\lambda\downarrow 0$), we have
if $\gamma<1$, then $\ \ \lim_{\lambda\downarrow 0}\frac{\sigma_{\varepsilon\nu}\big{(}1-\lambda v(-\lambda)\big{)}}{\alpha^{2}\big{(}1-\lambda p(-\lambda)\big{)}+\sigma_{\nu}^{2}(1-\lambda v(-\lambda))}=\frac{\gamma\sigma_{\varepsilon\nu}}{\alpha^{2}+\gamma\sigma^{2}_{\nu}}=\frac{\sigma_{\varepsilon\nu}}{\sigma^{2}_{\nu}}\ \frac{1}{\mu^{2}/k+1}\,,$
if $\gamma>1$, then $\ \ \lim_{\lambda\downarrow 0}\frac{\sigma_{\varepsilon\nu}\big{(}1-\lambda v(-\lambda)\big{)}}{\alpha^{2}\big{(}1-\lambda p(-\lambda)\big{)}+\sigma_{\nu}^{2}(1-\lambda v(-\lambda))}=\frac{\sigma_{\varepsilon\nu}}{\alpha^{2}+\sigma^{2}_{\nu}}=\frac{\sigma_{\varepsilon\nu}}{\sigma^{2}_{\nu}}\ \frac{1}{\mu^{2}/n+1}\,,$
using the definition of the concentration parameter $\mu^{2}\equiv n\alpha^{2}/\sigma_{\nu}^{2}\,$.
When we compare these results with the bias of OLS, which is
$\frac{\sigma_{\varepsilon\nu}}{\sigma^{2}_{\nu}}\ \frac{1}{\frac{\mu^{2}}{n}+1}\,,$
we find that the (ridgeless) 2SLS bias is never larger than the OLS bias (even if $\gamma>1$). The equivalence of the ridgeless 2SLS and OLS bias intuitively follows from the fact that the ridgeless 2SLS estimator interpolates the data in the first stage if $\gamma>1$, i.e. $x=Z\widehat{\pi}_{(\lambda\downarrow 0)}=\widehat{x}_{(\lambda\downarrow 0)}$. More interestingly, we can also show that increasing $\lambda$ reduces the bias relative to the (ridgeless) 2SLS bias (even if $\gamma>1$). To see this, we can rearrange the bias of 2SLS-Ridge (when we rule out the case $\lambda v(-\lambda)=1$, which occurs if and only if $\lambda\to\infty$)
$\frac{\sigma_{\varepsilon\nu}\big{(}1-\lambda v(-\lambda)\big{)}}{\alpha^{2}\big{(}1-\lambda p(-\lambda)\big{)}+\sigma_{\nu}^{2}(1-\lambda v(-\lambda))}=\frac{\sigma_{\varepsilon\nu}}{\sigma^{2}_{\nu}}\ \frac{1}{F\,a(-\lambda)+1}\,.$
$F\equiv\mu^{2}/k$ is the population first-stage $F$-statistic and $a(-\lambda)\equiv\gamma\frac{1-\lambda p(-\lambda)}{1-\lambda v(-\lambda)}$ can be interpreted as a Ridge amplifier of the first-stage $F$-statistic. We show in the next corollary, that this amplifier equals unity for the standard 2SLS (i.e., $\lambda=0$). Additionally, we show that its first derivative close to the ridgeless limit is positive, which suggests that there is a value of $\lambda$ for which the bias of 2SLS-Ridge is smaller compared to (ridgeless) 2SLS.
Corollary 4.1.2.
If $E[Y^{-1}]\geq 1$, $\gamma<1$ and $\lambda v(-\lambda)\neq 1$, then
$\lim_{\lambda\downarrow 0}a(-\lambda)=\lim_{\lambda\downarrow 0}\left(\frac{1}{p(-\lambda)}-\lambda\right)=1\,,$
$\lim_{\lambda\downarrow 0}a^{\prime}(-\lambda)=-\lim_{\lambda\downarrow 0}\left(\frac{p^{\prime}(-\lambda)}{p^{2}(-\lambda)}+1\right)\geq 0\,,$
with $-p^{\prime}(-\lambda)=m(-\lambda)-\lambda m^{\prime}(-\lambda)$ and $m^{\prime}(-\lambda)=\textrm{E}\left[\frac{1}{(Y+\lambda)^{2}}\right]\,$.
Proof. See Appendix D. ∎
Figure 1 depicts the bias approximation from Theorem 4.1 and the simulated bias of $\widetilde{\beta}_{\lambda}$ for different values of $\lambda$. In the left part of the figure we see the standard bias approximation for 2SLS if $\lambda\downarrow 0$ and $\gamma<1$, while in the right part we can observe that the ridgeless 2SLS bias equals the OLS bias if $\gamma>1\,$, which illustrates the results in Corollary 4.1.1. Independent of $\gamma$ we see that increasing $\lambda$ reduces the bias of 2SLS-Ridge. However, while increasing $\lambda$ reduces the bias of $\widetilde{\beta}_{\lambda}$, we cannot expect to fully get rid of it. We, therefore, propose a bias adjustment for the 2SLS-Ridge estimator in the following.
Figure 1: Bias approximation and simulation results for 2SLS-Ridge (average over 500 replications); dashed line shows RMT results from Theorem 4.1; solid grey line indicates OLS bias; AR-1 model $\Sigma_{ij}=0.5^{|i-j|}$
4.2 Bias Adjustment
The standard bias-adjusted 2SLS (aka Nagar’s estimator, 1959; Donald and Newey (2001); Chao and Swanson (2005)) has several representations
$\widehat{\beta}_{0}$ $=\frac{x^{\prime}P_{0}y-\frac{\gamma}{1-\gamma}x^{\prime}M_{0}y}{x^{\prime}P_{0}x-\frac{\gamma}{1-\gamma}x^{\prime}M_{0}x}$
$=\frac{(1-\gamma)x^{\prime}P_{0}y-\gamma x^{\prime}M_{0}y}{(1-\gamma)x^{\prime}P_{0}x-\gamma x^{\prime}M_{0}x}$
$=\frac{x^{\prime}P_{0}y-\gamma x^{\prime}y}{x^{\prime}P_{0}x-\gamma x^{\prime}x}\,.$
Analogue to the bias-adjusted OLS projection matrix, $P_{0}-\gamma I_{n}\,$, we propose a bias-adjusted Ridge smoother
$S_{\lambda}(s)=P_{\lambda}-sI_{n}\,.$
For the first term we know $\operatorname{tr}(P_{\lambda})/n\to_{a.s.}1-\lambda v(-\lambda)$. Therefore, considering
[…]
we know from the Silverstein equation that $\operatorname{tr}\big{(}S_{\lambda}(s_{1})\big{)}/n\to_{a.s.}0$. Unfortunately, the integral depends on unknown population values and cannot be derived in general. There are, however, several approaches to proceed. First, we may consider the case with $\lambda\downarrow 0$ and $\gamma<1$. Rearranging $s_{1}$ and using the dominated convergence theorem together with Lemma 3.3, we can derive that
[…]
because $\lim_{\lambda\downarrow 0}\lambda v(-\lambda)=1-\gamma$ and hence $\lim_{\lambda\downarrow 0}\lambda/\lambda v(-\lambda)=0$. Again invoking the Silverstein equation, gives us
$\operatorname{tr}\big{(}S_{0}(s_{1})\big{)}/n=\operatorname{tr}(P_{0}-\gamma I_{n})/n\to_{a.s.}0\,,$
which, for $\gamma<1$, can be used to construct the standard bias-adjusted 2SLS. Second, the availability of a closed-form expression for the Silverstein equation can be used to debias the 2SLS-Ridge estimator for values of $\lambda$ different from zero and/or values of $\gamma$ larger than unity. We illustrate this for isotropic instruments ($\Sigma=I_{k}$). When we consider
$s_{2}=\gamma\frac{v_{I}(-\lambda)}{1+v_{I}(-\lambda)}$
with the limit Stieltjes transform for such instruments depending only on $\gamma$ and $\lambda$
$\lambda v_{I}(-\lambda)=\frac{1-\gamma-\lambda+\sqrt{(1-\gamma-\lambda)^{2}+4\lambda}}{2}\,,$
then we know from the Silverstein equation that $\operatorname{tr}\big{(}S_{\lambda}(s_{2})\big{)}/n\to_{a.s.}0$.
Third, we can use an estimate of $v(-\lambda)$ to construct an $s$ for general matrices $\Sigma$ that fulfill Assumption 2. Consider
$s_{3}=1-\lambda\widehat{v}(-\lambda)$
with
$\widehat{v}(-\lambda)=\frac{1}{n}\operatorname{tr}\big{(}(ZZ^{\prime}/n+\lambda I_{n})^{-1}\big{)}\to_{a.s.}v(-\lambda)\,,$
and
$S_{\lambda}(s_{3})=\lambda\widehat{v}(-\lambda)P_{\lambda}-(1-\lambda\widehat{v}(-\lambda))M_{\lambda}$
such that by construction, $\operatorname{tr}\big{(}S_{\lambda}(s_{3})\big{)}/n=0\,$. We focus on $S_{\lambda}(s_{3})$, in the following simply referred to as $S_{\lambda}\,$, which gives us a bias-adjusted 2SLS-Ridge estimator that works for general values of $\lambda\,$,
$\widehat{\beta}_{\lambda}=\frac{x^{\prime}S_{\lambda}y}{x^{\prime}S_{\lambda}x}\,.$ (4.2)
The bias adjustment utilizes the fact that $P_{\lambda}$ and $M_{\lambda}$ act differently on the signal term and the noise term. By construction it successfully gets rid of the noise term while still keeping the signal alive. Theorem 4.2 shows that $\widehat{\beta}_{\lambda}$ estimates the causal effect consistently.
Theorem 4.2.
Under the conditions of Theorem 4.1, the bias of adjusted 2SLS-Ridge vanishes almost surely, i.e.,
$\widehat{\beta}_{\lambda}-\beta\to_{a.s.}0\,.$
Proof. See Appendix C. ∎
Using Lemma 3.3 we show in the next corollary that Theorem 4.2 also replicates the standard bias-adjusted 2SLS estimator as a special case (if $\gamma<1$). Additionally, it extends this estimator to ridgeless 2SLS (if $\gamma>1$).
Corollary 4.2.1.
For the ridgeless limit ($\lambda\downarrow 0$), we have
if γ<1, $\text{if $\gamma<1$, then}\ \lim$ ${}_{\lambda\downarrow 0}\,S_{\lambda}=(1-\gamma)P_{0}-\gamma M_{0}=P_{0}-\gamma I_{n}\,,$
if γ>1, $\text{if $\gamma>1$, then}\ \lim$ ${}_{\lambda\downarrow 0}\,S_{\lambda}=-M_{0}=P_{0}-I_{n}\,,$
with $P_{0}=Z(Z^{\prime}Z)^{-1}Z^{\prime}$ and $M_{0}=I_{n}-P_{0}\,$.
Figure 2 shows that $\widehat{\beta}_{\lambda}$ is (by construction) unbiased for all values of $\lambda$ (including $\lambda\downarrow 0$). The bias adjustment works also if $\gamma>1\,$ (given that $\lambda\neq 0$).
Figure 2: Bias approximation and simulation results for bias-adjusted 2SLS-Ridge (average over 500 replications); dashed line shows RMT results from Theorem 4.2; solid grey line indicates OLS bias; AR-1 model $\Sigma_{ij}=0.5^{|i-j|}$
It is well known that the standard bias-adjusted 2SLS estimator (i.e., $\lambda=0$ and $\gamma<1$) can perform poorly in finite samples. A reason for this is that $\gamma$ in Corollary 4.2.1 can be interpreted as costs for the bias adjustment, which reduce the signal to $\alpha^{2}(1-\gamma)$. By inspecting the limit of the denominator of bias-adjusted 2SLS-Ridge derived in the proof of Theorem 4.2, we see two negative terms:
$\frac{x^{\prime}S_{\lambda}x}{n}\to_{a.s.}\alpha^{2}\big{(}1-(\lambda+\gamma)p(-\lambda)\big{)}\equiv\alpha^{2}f(-\lambda)\,.$
The first term, which is also in $\widetilde{\beta}_{\lambda}$, stems from using Ridge ($-\lambda p(-\lambda)$) and the second term stems from the bias adjustment ($-\gamma p(-\lambda)$). In both cases $p(-\lambda)$ can be interpreted as a price per unit, which is unity if $\lambda=0$ and declines with increasing $\lambda$. This decline leads to a trade-off that we can use to make bias adjustments less costly (with respect to the signal term).
The next corollary discusses this trade-off close to the ridgeless limit and shows under what condititons we can expect to see a larger signal term if $\lambda$ increases. Compared with standard bias-adjusted 2SLS, the trade-off turns out to be positive if $\gamma\textrm{E}\left[Y^{-1}\right]\geq 1$. This condition simplifies to $\gamma>1/2\,$ for isotropic instruments (i.e., $E\left[T^{-1}\right]=1$) because from the Marchenko-Pastur equation we know $E(Y^{-1})=E(T^{-1})/(1-\gamma)$.
Corollary 4.2.2.
If $\gamma E[Y^{-1}]\geq 1$, $\gamma<1$ and $\lambda v(-\lambda)\neq 1$, then
$\lim_{\lambda\downarrow 0}f(-\lambda)=\lim_{\lambda\downarrow 0}\big{(}1-(\lambda+\gamma)p(-\lambda)\big{)}=1-\gamma$
$\lim_{\lambda\downarrow 0}f^{\prime}(-\lambda)=-\lim_{\lambda\downarrow 0}\big{(}(\lambda+\gamma)p^{\prime}(-\lambda)+p(-\lambda)\big{)}\geq 0\,.$
Proof. See Appendix D. ∎
Figure 3: Signal approximation and simulation results for $x^{\prime}S_{\lambda}x/n$ (average over 500 replications) and $\alpha^{2}f(-\lambda)$ (dashed line); solid grey line indicates signal from standard bias-adjusted 2SLS, i.e. $\alpha^{2}(1-\gamma)$ if $\gamma<1$ and zero otherwise; $\Sigma=I_{k}$
Figure 3 illustrates the trade-off described in Corollary 4.2.2 for different values of $\gamma$ and $\lambda$. If $\gamma>1/2$, the signal can be improved compared to standard bias-adjusted 2SLS. If $\gamma>1$, we know for standard bias-adjusted 2SLS that $\lim_{\lambda\downarrow 0}f(-\lambda)=0$ (using Lemma 3.3) implying that standard bias-adjusted 2SLS is not applicable. However, a positive value of $\lambda$ can reestablish the signal, which enables us to identify the causal effect even if $\gamma>1$. We see similar patterns also for the AR-1 covariance matrix (see Figure 6 in Appendix F).
4.3 Variance and Cross Validation
For inference we consider the Bekker (1994) standard errors. Let $\widehat{\varepsilon}=y-x\widehat{\beta}_{\lambda}$, $\widehat{\sigma}_{\varepsilon}^{2}=\widehat{\varepsilon}^{\prime}\widehat{\varepsilon}/n$, and $\widetilde{x}=x+\widehat{\varepsilon}\,(\widehat{\varepsilon}^{\prime}x)/(\widehat{\varepsilon}^{\prime}\widehat{\varepsilon})$. Inspired by Hansen et al. (2008), we can estimate the variance of the standard bias-adjusted 2SLS estimator by
$\widehat{\textrm{Var}}(\widehat{\beta}_{0})=\widehat{\sigma}_{\varepsilon}^{2}\,\frac{x^{\prime}S_{0}S_{0}\widetilde{x}}{(x^{\prime}S_{0}x)^{2}}\,.$
Following this structure, we propose to estimate the variance of the bias-adjusted 2SLS-Ridge estimator by
$\widehat{\textrm{Var}}(\widehat{\beta}_{\lambda})=\widehat{\sigma}_{\varepsilon}^{2}\,\frac{x^{\prime}S_{\lambda}S_{\lambda}\widetilde{x}}{(x^{\prime}S_{\lambda}x)^{2}}\,.$
The corresponding asymptotic variance of $\widehat{\beta}_{\lambda}$ is (see Appendix E for a derivation)
$\widehat{\textrm{Var}}(\widehat{\beta}_{\lambda})\to_{a.s.}\frac{1}{n}\,\left(\frac{\sigma_{\varepsilon}^{2}}{\alpha^{2}}\frac{(1-\gamma p)f-\lambda p(1-\gamma p)-\lambda^{2}p^{\prime}}{f^{2}}+\frac{\left(\sigma^{2}_{\varepsilon}\sigma_{v}^{2}+\sigma_{\varepsilon\nu}^{2}\right)}{\alpha^{4}}\frac{\big{[}\lambda^{2}v^{\prime}-\lambda^{2}v^{2}\big{]}}{f^{2}}\right)\,,$
which for $\lambda\downarrow 0$ coincides with the well-known asymptotic variance of the standard bias-adjusted 2SLS estimator under normality of the errors (compare, e.g., Bekker, 1994; Anatolyev, 2013; Kolesár et al., 2015; Anatolyev, 2019)
$\textrm{Var}(\widehat{\beta}_{0})$ $=\frac{1}{n}\left(\frac{\sigma_{\varepsilon}^{2}}{\alpha^{2}}+\frac{\gamma}{1-\gamma}\frac{\sigma_{\varepsilon}^{2}\sigma_{\nu}^{2}+\sigma_{\varepsilon\nu}^{2}}{\alpha^{4}}\right)\,.$
Figure 4 compares the estimated Bekker-type variance for the bias-adjusted 2SLS-Ridge estimator with the observed variation of $\widehat{\beta}_{\lambda}$ in the simulation. We further add the asymptotic variance of the standard bias-adjusted 2SLS and the variance of the Liml estimator as benchmarks (for $\gamma<1$ only).
Figure 4: Observed variation and estimated variance of $\widehat{\beta}_{\lambda}$ (average over 500 replications) and asymptotic variance (dashed line). The upper and lower horizontal lines depict the asymptotic variance of the standard bias-adjusted 2SLS and Liml, respectively; top: $\Sigma=I_{k}$; bottom: AR-1 model $\Sigma_{ij}=0.5^{|i-j|}$
Cross validation is a method to find an optimally tuned $\lambda$ by evaluating the observed bias-variance trade-off. Because $\widehat{\beta}_{\lambda}$ is already consistent for all values of $\lambda$, we use cross validation to find the optimal value of $\lambda$ that minimizes the variance of $\widehat{\beta}_{\lambda}$. To obtain this, we use the following criterion for cross validation
$CV(\lambda)=\ln(x^{\prime}S_{\lambda}S_{\lambda}\widetilde{x})-2\ln(x^{\prime}S_{\lambda}x)\,.$
In Figure 5 we illustrate the behavior of bias-adjusted 2SLS-Ridge for the optimal $\lambda$ derived from minimizing $CV(\lambda)$. For comparison, we also depict the results from (ridgeless) 2SLS/bias-adjusted 2SLS, and Liml (if $\gamma<1$). All bias-adjusted estimators succesfully eliminate the bias that we observe for (ridgeless) 2SLS. However, we see that our bias-adjusted 2SLS-Ridge can be more efficient than Liml and, additionally, more efficient than a ridgeless version ($\lambda\downarrow 0$) of the standard bias-adjusted 2SLS estimator if $\gamma=1.25$.
Figure 5: Density plots of estimates (average over 1,000 replications); AR-1 model $\Sigma_{ij}=0.5^{|i-j|}\,$; top: $F=1$; middle: $F=2$; bottom: $F=5$
A Appendix: Proof of Lemmas
Proof of Lemma 3.1
Proof.
$I_{n}$ $=$ $(ZZ^{\prime}/n+\lambda I_{n})(ZZ^{\prime}/n+\lambda I_{n})^{-1}$
$=$ $(ZZ^{\prime}/n)(ZZ^{\prime}/n+\lambda I_{n})^{-1}+\lambda(ZZ^{\prime}/n+\lambda I_{n})^{-1}$
$=$ $Z(Z^{\prime}Z/n+\lambda I_{k})^{-1}Z^{\prime}/n+\lambda(ZZ^{\prime}/n+\lambda I_{n})^{-1}$
The last equation, which proves the Lemma, uses the so-called “kernel trick”, i.e. $(Z^{\prime}Z/n+\lambda I_{k})^{-1}Z^{\prime}=Z^{\prime}(ZZ^{\prime}/n+\lambda I_{n})^{-1}\,$. ∎
Proof of Lemma 3.2
Proof. This identity can be found in many textbooks (see e.g. Lemma 2.4 in Couillet and Liao, 2022). It follows because the ESDs of $\widehat{\Sigma}$ and $\underline{\widehat{\Sigma}}$ have the same nonzero eigenvalues and differ only by $|n-k|$ zero eigenvalues for the larger matrix, that is
$v(-\lambda)=\gamma m(-\lambda)+(1-\gamma)\frac{1}{\lambda}\,.$
Multiplying by $\lambda$ and rearranging leads to the claimed formula. ∎
Proof of Lemma 3.3
Proof. The proof follows from Lemma 2.3 in Dobriban and Wager (2018a). They use the dominated convergence theorem to swap limits and expectations, which allows us to take the limit as $\lambda\downarrow 0$ inside the expectation. If $\gamma<1$, we know that $Y$ is bounded away from 0 ($Y>c>0$ for some $c$) and the dominated convergence theorem is hence applicable to $m(-\lambda)$. If $\gamma>1$, we know that $\underline{Y}$ is bounded away from 0 ($\underline{Y}>c>0$ for some $c$) and the dominated convergence theorem is applicable to $v(-\lambda)$. We first consider the case with $\gamma<1$, then
$\lim_{\lambda\downarrow 0}\lambda m(-\lambda)=\lim_{\lambda\downarrow 0}\textrm{E}\left[\frac{\lambda}{Y+\lambda}\right]=\textrm{E}\left[\frac{0}{Y}\right]=0\,,$
which proves 1b). 1a) follows immediately from this result using Lemma 3.2. Similarly, if $\gamma>1$, then
$\lim_{\lambda\downarrow 0}\lambda v(-\lambda)=\lim_{\lambda\downarrow 0}\textrm{E}\left[\frac{\lambda}{\underline{Y}+\lambda}\right]=\textrm{E}\left[\frac{0}{\underline{Y}}\right]=0\,,$
which proves 2a). 2b) follows immediately from this result using Lemma 3.2. ∎
B Appendix: Proof of Theorem 4.1
Proof. By definition of the linear model (2.1) we can write
$\widetilde{\beta}_{\lambda}=\beta+\frac{x^{\prime}P_{\lambda}\varepsilon}{x^{\prime}P_{\lambda}x}\,.$
To establish convergence of the remainder term, we argue that the quadratic terms involving $P_{\lambda}$ concentrate around their means and then use results from RMT to find the limits of those means. We begin with the denominator.
Following Hastie et al. (2022a), let $x=Z\pi+\nu$ and $c_{n}=\sqrt{k}(\sigma_{\nu}/\alpha)$ and observe that
$\frac{x^{\prime}P_{\lambda}x}{n}$ $=(\pi,\nu)^{\prime}\left(\frac{1}{n}\begin{bmatrix}Z\\ I_{n}\end{bmatrix}^{\prime}P_{\lambda}\begin{bmatrix}Z\\ I_{n}\end{bmatrix}\right)(\pi,\nu)$
$=\underbrace{(\pi,\nu/c_{n})^{\prime}}_{\delta^{\prime}}\underbrace{\left(\frac{1}{n}\begin{bmatrix}Z\\ c_{n}I_{n}\end{bmatrix}^{\prime}P_{\lambda}\begin{bmatrix}Z\\ c_{n}I_{n}\end{bmatrix}\right)}_{A}\underbrace{(\pi,\nu/c_{n})}_{\delta}.$
Because $\delta$ and $A$ are independent and $\delta$ has independent entries with mean zero and common variance $\alpha^{2}/k$, we can use the almost sure convergence of quadratic forms, from Lemma C.3 in Dobriban and Wager (2018b), which is adapted from Lemma B.26 in Bai and Silverstein (2010). This result asserts that, almost surely,
$\delta^{\prime}A\delta-\left(\frac{\alpha^{2}}{k}\right)\operatorname{tr}{(A)}\to_{a.s.}0\,.$
Now examine
$\frac{\alpha^{2}}{k}\operatorname{tr}{(A)}$ $=\frac{\alpha^{2}}{k}\operatorname{tr}\left(P_{\lambda}\left(\frac{ZZ^{\prime}}{n}+\frac{c_{n}^{2}}{n}I_{n}\right)\right)$
$=\frac{\alpha^{2}}{k}\operatorname{tr}\left(Z^{\prime}P_{\lambda}Z/n\right)+\frac{\sigma_{\nu}^{2}}{n}\operatorname{tr}\left(P_{\lambda}\right)$
$=\underbrace{\frac{\alpha^{2}}{k}\operatorname{tr}\left(Z^{\prime}Z/n\right)}_{a}-\underbrace{\frac{\alpha^{2}\lambda}{k}\operatorname{tr}\big{(}Z^{\prime}(ZZ^{\prime}/n+\lambda I_{n})^{-1}Z/n\big{)}}_{b}+\underbrace{\frac{\sigma_{\nu}^{2}}{n}\operatorname{tr}\left(P_{\lambda}\right)}_{c}\,.$
where the first and second equality follow from cyclic rotations within the trace operator and the third equality uses Lemma 3.1. In the following we consider $a,b$ and $c$ separately. For convenience, we use the Stieltjes transform $m(-\lambda)$ for all terms related with the identification strength $\alpha^{2}$ and the companion Stieltjes transform $v(-\lambda)$ for all terms related with $\sigma^{2}_{\nu}$ or $\sigma_{\varepsilon\nu}$. It is known that
$a=\frac{\alpha^{2}}{k}\operatorname{tr}\left(Z^{\prime}Z/n\right)\to_{a.s.}\alpha^{2}\,.$
Moreover,
$b$ $=\frac{\alpha^{2}\lambda}{k}\operatorname{tr}\big{(}Z^{\prime}(ZZ^{\prime}/n+\lambda I_{n})^{-1}Z/n\big{)}$
$=\frac{\alpha^{2}\lambda}{k}\operatorname{tr}\big{(}(Z^{\prime}Z/n)(Z^{\prime}Z/n+\lambda I_{k})^{-1}\big{)}$
$=\frac{\alpha^{2}\lambda}{k}\operatorname{tr}\big{(}I_{k}-\lambda(Z^{\prime}Z/n+\lambda I_{k})^{-1}\big{)}\to_{a.s.}\alpha^{2}\lambda\left(1-\lambda m(-\lambda)\right)\,,$
where the second equality uses the kernel trick again and the third equality follows from the relation $Z^{\prime}Z/n=(Z^{\prime}Z/n+\lambda I_{k})-\lambda I_{k}\,$.
Finally, by similar calculations, we have
$c$ $=\frac{\sigma_{\nu}^{2}}{n}\operatorname{tr}\big{(}Z(Z^{\prime}Z/n+\lambda I_{k})^{-1}Z^{\prime}/n\big{)}$
$=\frac{\sigma_{\nu}^{2}}{n}\operatorname{tr}\big{(}I_{n}-\lambda(ZZ^{\prime}/n+\lambda I_{n})^{-1}\big{)}$
$\to_{a.s.}\sigma_{\nu}^{2}(1-\lambda v(-\lambda))\,.$
Hence, we have shown that, almost surely,
$\frac{x^{\prime}P_{\lambda}x}{n}\to_{a.s.}\alpha^{2}\big{(}1-\lambda(1-\lambda m(-\lambda))\big{)}+\sigma_{\nu}^{2}(1-\lambda v(-\lambda))\,.$
For the numerator we get
$\frac{x^{\prime}P_{\lambda}\varepsilon}{n}$ $=(Z\pi+\nu)^{\prime}P_{\lambda}\varepsilon/n$
$=\pi^{\prime}Z^{\prime}P_{\lambda}\varepsilon/n+\frac{1}{n}\operatorname{tr}\big{(}\nu^{\prime}(I_{n}-\lambda(ZZ^{\prime}/n+\lambda I_{n})^{-1})\varepsilon\big{)}$
$\to_{a.s.}\sigma_{\varepsilon\nu}\big{(}1-\lambda v(-\lambda)\big{)}\,.$
where the second equality uses Lemma 3.1. Hence, by the continuous mapping theorem, we have
$\widetilde{\beta}_{\lambda}-\beta\to_{a.s.}\frac{\sigma_{\varepsilon\nu}\big{(}1-\lambda v(-\lambda)\big{)}}{\alpha^{2}\big{(}1-\lambda(1-\lambda m(-\lambda))\big{)}+\sigma_{\nu}^{2}(1-\lambda v(-\lambda))}\,.$
C Appendix: Proof of Theorem 4.2
Proof. By definition of the linear model (2.1) we can write
$\widehat{\beta}_{\lambda}=\beta+\frac{x^{\prime}S_{\lambda}\varepsilon}{x^{\prime}S_{\lambda}x}\,.$
Recall the definition of $S_{\lambda}=\lambda\widehat{v}(-\lambda)P_{\lambda}-(1-\lambda\widehat{v}(-\lambda))M_{\lambda}\,$. From Theorem 4.1 we know for the first term of the denominator that
$\lambda\widehat{v}(-\lambda)\frac{x^{\prime}P_{\lambda}x}{n}\to_{a.s.}\alpha^{2}\big{(}1-\gamma p(-\lambda)\big{)}\big{(}1-\lambda p(-\lambda)\big{)}+\sigma_{\nu}^{2}\lambda v(-\lambda)(1-\lambda v(-\lambda))\,.$
where we use Lemma 3.2 and the definition $p(-\lambda)=1-\lambda m(-\lambda)$ for clarity.
Using similar calculations as in Theorem 4.1, we get for the second term
$\frac{x^{\prime}M_{\lambda}x}{n}$ $=(\pi,\nu)^{\prime}\left(\frac{1}{n}\begin{bmatrix}Z\\ I_{n}\end{bmatrix}^{\prime}(I_{n}-P_{\lambda})\begin{bmatrix}Z\\ I_{n}\end{bmatrix}\right)(\pi,\nu)$
$=\underbrace{(\pi,\nu/c_{n})^{\prime}}_{\delta^{\prime}}\underbrace{\left(\frac{1}{n}\begin{bmatrix}Z\\ c_{n}I_{n}\end{bmatrix}^{\prime}(I_{n}-P_{\lambda})\begin{bmatrix}Z\\ c_{n}I_{n}\end{bmatrix}\right)}_{A}\underbrace{(\pi,\nu/c_{n})}_{\delta}.$
Again because $\delta$ and $A$ are independent and $\delta$ has independent entries with mean zero and common variance $\alpha^{2}/k$, we can use the almost sure convergence of quadratic forms, from Lemma C.3 in Dobriban and Wager (2018b), which is adapted from Lemma B.26 in Bai and Silverstein (2010). Therefore, we have
$\frac{\alpha^{2}}{k}\operatorname{tr}{(A)}$ $=\frac{\alpha^{2}}{k}\operatorname{tr}\left(M_{\lambda}\left(\frac{ZZ^{\prime}}{n}+\frac{c_{n}^{2}}{n}I_{n}\right)\right)$
$=\frac{\alpha^{2}}{k}\operatorname{tr}\left(Z^{\prime}M_{\lambda}Z/n\right)+\frac{\sigma_{\nu}^{2}}{n}\operatorname{tr}\left(M_{\lambda}\right)$
$=\underbrace{\frac{\alpha^{2}\lambda}{k}\operatorname{tr}\big{(}Z^{\prime}(ZZ^{\prime}/n+\lambda I_{n})^{-1}Z/n\big{)}}_{a}+\underbrace{\frac{\sigma_{\nu}^{2}}{n}\operatorname{tr}\left(M_{\lambda}\right)}_{b}\,,$
where the first and second equality follow from cyclic rotations within the trace operator and the third equality uses Lemma 3.1. Again, we have
$a$ $=\frac{\alpha^{2}\lambda}{k}\operatorname{tr}\big{(}Z^{\prime}(ZZ^{\prime}/n+\lambda I_{n})^{-1}Z/n\big{)}$
$=\frac{\alpha^{2}\lambda}{k}\operatorname{tr}\big{(}(Z^{\prime}Z/n)(Z^{\prime}Z/n+\lambda I_{k})^{-1}\big{)}$
$=\frac{\alpha^{2}\lambda}{k}\operatorname{tr}\big{(}I_{k}-\lambda(Z^{\prime}Z/n+\lambda I_{k})^{-1}\big{)}\to_{a.s.}\alpha^{2}\lambda p(-\lambda)\,.$
where the second equality uses the kernel trick again and the third equality follows from the relation $Z^{\prime}Z/n=(Z^{\prime}Z/n+\lambda I_{k})-\lambda I_{k}\,$.
Finally, by similar calculations, we have
$b$ $=\frac{\sigma_{\nu}^{2}}{n}\operatorname{tr}\left(M_{\lambda}\right)$
$=\sigma_{\nu}^{2}\lambda\operatorname{tr}\left(\frac{1}{n}(ZZ^{\prime}/n+\lambda I_{n})^{-1}\right)$
$\to_{a.s.}\sigma_{\nu}^{2}\lambda v(-\lambda)\,,$
where the second equality follows from Lemma 3.1. Using Lemma 3.2, we can show that, almost surely,
(1(1-( 1 - $\lambda\widehat{v}(-\lambda))\frac{x^{\prime}M_{\lambda}x}{n}$
$\to_{a.s.}\alpha^{2}\gamma\lambda p^{2}(-\lambda)+\sigma_{\nu}^{2}(1-\lambda v(-\lambda))\lambda v(-\lambda)\,.$
Therefore, the denominator of $\widehat{\beta}$ converges almost surely to
$\frac{x^{\prime}S_{\lambda}x}{n}\to_{a.s.}\alpha^{2}\big{(}1-(\lambda+\gamma)p(-\lambda)\big{)}\,,$
where the terms related to $\sigma^{2}_{\nu}$ cancel out. For the numerator we first need
$\frac{x^{\prime}M_{\lambda}\varepsilon}{n}$ $=(Z\pi+\nu)^{\prime}M_{\lambda}\varepsilon/n$
$=\frac{\lambda}{n}\operatorname{tr}\big{(}(Z\pi+\nu)^{\prime}(ZZ^{\prime}/n+\lambda I_{n})^{-1}\varepsilon\big{)}$
$\to_{a.s.}\sigma_{\varepsilon\nu}\lambda v(-\lambda)\,,$
and, combined with the results from Theorem 4.1, eventually get
$\frac{x^{\prime}S_{\lambda}\varepsilon}{n}$ $\to_{a.s.}\sigma_{\varepsilon\nu}\lambda v(-\lambda)\big{(}1-\lambda v(-\lambda)\big{)}-\big{(}1-\lambda v(-\lambda)\big{)}\sigma_{\varepsilon\nu}\lambda v(-\lambda)=0\,.$
Hence, by the continuous mapping theorem, we finally have
$\widehat{\beta}_{\lambda}-\beta\to_{a.s.}0\,.$
D Appendix: Proof of Corollaries
To simplify the illustration, we use the following notation in the proofs of the Corollaries: $m^{\prime}(-\lambda)=\textrm{E}\left[\frac{1}{(Y+\lambda)^{2}}\right]$ and $-p^{\prime}(-\lambda)=m(-\lambda)-\lambda m^{\prime}(-\lambda)\,$.
Proof of Corollary 4.1.2
Proof. The first claim, $\lim_{\lambda\downarrow 0}a(-\lambda)=1$, follows immediately from Lemma 3.3. For the second claim, note that
$a(-\lambda)=\gamma\frac{1-\lambda p(-\lambda)}{1-\lambda v(-\lambda)}=\frac{1-\lambda p(-\lambda)}{p(-\lambda)}=\frac{1}{p(-\lambda)}-\lambda$
which follows from Lemma 3.2. For the first derivative we see
$a^{\prime}(-\lambda)=-\frac{p^{\prime}(-\lambda)}{p^{2}(-\lambda)}-1\,,$
which, for $v(-\lambda)\neq 1$ or equivalently $m(-\lambda)\neq 1$, is positive if
$m(-\lambda)-\lambda m^{\prime}(-\lambda)-1+2\lambda m(-\lambda)-\lambda^{2}m^{2}(-\lambda)\geq 0\,,$
or, equivalently,
$\underbrace{m(-\lambda)-\lambda m^{\prime}(-\lambda)-1+2\lambda m(-\lambda)-\lambda^{2}m^{\prime}(-\lambda)}_{A}+\underbrace{\lambda^{2}m^{\prime}(-\lambda)-\lambda^{2}m^{2}(-\lambda)}_{B}\geq 0\,.$
Note that
$B$ $=\textrm{E}\left[\left(\frac{\lambda}{Y+\lambda}\right)^{2}\right]-\textrm{E}\left[\frac{\lambda}{Y+\lambda}\right]^{2}\geq 0$
is a variance term and hence positive, and
$A$ $=\textrm{E}\left[\frac{1}{Y+\lambda}\right]-\textrm{E}\left[\frac{\lambda+\lambda^{2}}{(Y+\lambda)^{2}}\right]+\textrm{E}\left[\frac{2\lambda}{Y+\lambda}\right]-1$
$=\textrm{E}\left[\frac{(1+2\lambda)(Y+\lambda)-\lambda-\lambda^{2}-(Y^{2}+2\lambda Y+\lambda^{2})}{(Y+\lambda)^{2}}\right]$
$=\textrm{E}\left[\frac{Y-Y^{2}}{(Y+\lambda)^{2}}\right].$
Moreover, using the dominated convergence theorem we get
$\lim_{\lambda\downarrow 0}\textrm{E}\left[\frac{Y-Y^{2}}{(Y+\lambda)^{2}}\right]=E[Y^{-1}]-1\geq 0\,,$
which proves the claim. ∎
Proof of Corollary 4.2.2
The first claim, $\lim_{\lambda\downarrow 0}f(\lambda)=1-\gamma$, follows immediately from Lemma 3.3. For the second claim, note that
$f^{\prime}(-\lambda)$ $=-\big{(}\lambda+\gamma\big{)}p^{\prime}(-\lambda)-p(-\lambda)\,$
Therefore, we get
$f^{\prime}(-\lambda)=\textrm{E}\left[\frac{(\lambda+\gamma)Y}{(Y+\lambda)^{2}}\right]-\textrm{E}\left[\frac{Y}{Y+\lambda}\right]=\gamma\textrm{E}\left[\frac{Y}{(Y+\lambda)^{2}}\right]-\textrm{E}\left[\frac{Y^{2}}{(Y+\lambda)^{2}}\right]\,,$
Moreover, using the dominated convergence theorem we get
$\lim_{\lambda\downarrow 0}\left(\gamma\textrm{E}\left[\frac{Y}{(Y+\lambda)^{2}}\right]-\textrm{E}\left[\frac{Y^{2}}{(Y+\lambda)^{2}}\right]\right)=\gamma\textrm{E}\left[Y^{-1}\right]-1\geq 0\,,$
which proves the claim.
E Appendix: Asymptotic Variance of $\widehat{\beta}_{\lambda}$
From Theorem 4.2 we have $x^{\prime}S_{\lambda}x/n\to_{a.s.}\alpha^{2}f\,$, where we suppress the $(-\lambda)$ in the following to simplify the notation. Moreover,
$\frac{x^{\prime}S_{\lambda}S_{\lambda}x}{n}$ $=(\lambda\widehat{v})^{2}\,\frac{x^{\prime}P_{\lambda}P_{\lambda}x}{n}+(1-\lambda\widehat{v})^{2}\,\frac{x^{\prime}M_{\lambda}M_{\lambda}x}{n}-2\lambda\widehat{v}(1-\lambda\widehat{v})\,\frac{x^{\prime}P_{\lambda}M_{\lambda}x}{n}$
$\to_{a.s.}\alpha^{2}\big{[}(1-\gamma p)f-\lambda p(1-\gamma p)-\lambda^{2}p^{\prime}\big{]}+\sigma_{v}^{2}\big{[}\lambda^{2}v^{\prime}-\lambda^{2}v^{2}\big{]}$
First, consider the numerator of $\widehat{\textrm{Var}}(\widehat{\beta}_{\lambda})$,
$\frac{x^{\prime}S_{\lambda}S_{\lambda}\widetilde{x}}{n}$ $=(\lambda\widehat{v})^{2}\,\frac{x^{\prime}P_{\lambda}P_{\lambda}\widetilde{x}}{n}+(1-\lambda\widehat{v})^{2}\,\frac{x^{\prime}M_{\lambda}M_{\lambda}\widetilde{x}}{n}-2\lambda\widehat{v}(1-\lambda\widehat{v})\,\frac{x^{\prime}P_{\lambda}M_{\lambda}\widetilde{x}}{n}$
$=\frac{x^{\prime}S_{\lambda}S_{\lambda}x}{n}+\frac{1}{n}\big{(}x^{\prime}P_{\lambda}P_{\lambda}\widehat{\varepsilon}-2(1-\lambda\widehat{v})x^{\prime}P_{\lambda}\widehat{\varepsilon}+(1-\lambda\widehat{v})^{2}x^{\prime}\widehat{\varepsilon}\,\big{)}\frac{\widehat{\varepsilon}^{\prime}x}{\widehat{\varepsilon}^{\prime}\widehat{\varepsilon}}$
$=\frac{x^{\prime}S_{\lambda}S_{\lambda}x}{n}+\frac{1}{n}\big{(}x^{\prime}P_{\lambda}P_{\lambda}\widehat{\varepsilon}-(1-\lambda\widehat{v})^{2}x^{\prime}\widehat{\varepsilon}\,\big{)}\frac{\widehat{\varepsilon}^{\prime}x}{\widehat{\varepsilon}^{\prime}\widehat{\varepsilon}}$
$=\frac{x^{\prime}S_{\lambda}S_{\lambda}x}{n}+\big{(}\lambda^{2}\widehat{v}^{\prime}-\lambda^{2}\widehat{v}^{2}\big{)}\frac{x^{\prime}\widehat{\varepsilon}}{n}\,\frac{\widehat{\varepsilon}^{\prime}x}{\widehat{\varepsilon}^{\prime}\widehat{\varepsilon}}$
$\to_{a.s.}\alpha^{2}\big{[}(1-\gamma p)f-\lambda p(1-\gamma p)-\lambda^{2}p^{\prime}\big{]}+\left(\sigma_{v}^{2}+\frac{\sigma_{\varepsilon\nu}^{2}}{\sigma^{2}_{\varepsilon}}\right)\big{[}\lambda^{2}v^{\prime}-\lambda^{2}v^{2}\big{]}\,.$
where we use $S_{\lambda}S_{\lambda}=P_{\lambda}P_{\lambda}-2(1-\lambda\widehat{v})P_{\lambda}+(1-\lambda\widehat{v})^{2}I_{n}$ in the second equality, $x^{\prime}P_{\lambda}\widehat{\varepsilon}=(1-\lambda\widehat{v})x^{\prime}\widehat{\varepsilon}$ in the third equality, and $x^{\prime}P_{\lambda}P_{\lambda}\widehat{\varepsilon}=(1-2\lambda\widehat{v}+\lambda^{2}\widehat{v}^{\prime})x^{\prime}\widehat{\varepsilon}$ in the fourth equality. Eventually, we get
$\widehat{\textrm{Var}}(\widehat{\beta}_{\lambda})\to_{a.s.}\frac{\sigma_{\varepsilon}^{2}}{n}\,\frac{\alpha^{2}\big{[}(1-\gamma p)f-\lambda p(1-\gamma p)-\lambda^{2}p^{\prime}\big{]}+\left(\sigma_{v}^{2}+\frac{\sigma_{\varepsilon\nu}^{2}}{\sigma^{2}_{\varepsilon}}\right)\big{[}\lambda^{2}v^{\prime}-\lambda^{2}v^{2}\big{]}}{\alpha^{4}f^{2}}\,.$
For $\lambda=0$ and $\gamma<1\,$, we get the Bekker asymptotic variance as special case because then $\lim_{\lambda\downarrow 0}p(-\lambda)=1$, $\lim_{\lambda\downarrow 0}p^{\prime}(-\lambda)=\textrm{E}[Y^{-1}]$, $\lim_{\lambda\downarrow 0}f(-\lambda)=1-\gamma$ and
$\lim_{\lambda\downarrow 0}\lambda^{2}v^{\prime}(-\lambda)-\lambda^{2}v^{2}(-\lambda)$ $=\lim_{\lambda\downarrow 0}\lambda^{2}v^{2}(-\lambda)\big{(}\widehat{v}^{\prime}(-\lambda)/\widehat{v}^{2}(-\lambda)-1\big{)}$
$=(1-\gamma)^{2}\left(\frac{1}{1-\gamma}-1\right)=\gamma(1-\gamma)\,.$
Therefore,
$\widehat{\textrm{Var}}(\widehat{\beta}_{0})\to_{a.s.}\frac{1}{n}\left(\frac{\sigma_{\varepsilon}^{2}}{\alpha^{2}}+\frac{\gamma}{1-\gamma}\frac{\sigma_{\varepsilon}^{2}\sigma_{\nu}^{2}+\sigma_{\varepsilon\nu}^{2}}{\alpha^{4}}\right)=\textrm{Var}(\widehat{\beta}_{0})\,.$
F Appendix: Additional Simulation Results
Figure 6: Signal approximation and simulation results for $x^{\prime}S_{\lambda}x/n$ (average over 500 replications) and $\alpha^{2}f(-\lambda)$ (dashed line); solid grey line indicates signal from standard bias-adjusted 2SLS, i.e. $\alpha^{2}(1-\gamma)$ if $\gamma<1$ and zero otherwise; AR-1 model $\Sigma_{ij}=0.5^{|i-j|}$